A 1300 kg sports car takes 3.88 seconds to reach a speed 27.7 mls. IT consumes 1.55*10^6J of chemical energy to do this. The average power output by the car's engine as mechanical energy is a.bc*10^d W. the value of a,b,c and d are .....
1
Expert's answer
2020-06-15T10:22:20-0400

Power is energy divided by time:


"P=\frac{E}{t},\\\space\\
P=\frac{1.55\cdot10^6}{3.88}=3.99\cdot10^5\text{ W}."

The value of a, b, c, and d are 3, 9, 9, 5 respectively.